Kerosene leaks out a hole 0.50 in.in diameterat thebottom of a tank at a rate of 10gal/min. How high is the kerosene in the tank?


1
Expert's answer
2021-08-11T09:54:12-0400

Given:

"d=0.50\:\rm in=1.27\: cm"

"\frac{\Delta V}{\Delta t}=10\:\rm gal/min=6.31*10^{-4}\: m^3/s"

The flow rate

"\frac{\Delta V}{\Delta t}=\frac{Av\Delta t}{\Delta t}=Av=\pi d^2/4*v"

The speed of a flow

"v=\sqrt{2gH}"

Therefore,

"H=\frac{(4\frac{\Delta V}{\Delta t}/\pi d^2)^2}{2g}"

"=\frac{(4*6.31*10^{-4}/3.14*(1.27*10^{-2})^2)^2}{2*9.8}=1.27\:\rm m"